PV of Annuity Due

1) if 10,000 must be withdrawn at the beginning of each 3months period for the next 3 years, how much must be deposited now, at 6% compounded quarterly, to yield the annuity payment needed?

2) billy wants 5000 at the end of each 3 month period for the next 6 years. If billy banks is paying 8% interest compounded quarterly, how much must she deposit if she wants the payments at the beginning of each 3 month period?


1
Expert's answer
2021-05-20T17:19:25-0400

1) n=3*4=12, pmt= 10000 , rate "r=\dfrac{0.06}{4}=0.015"


In Excel: =PV(0.015,12,10000,0,1)

      =62337.12


2) n=6*4=24,pmt=5000, rate "r=\dfrac{0.08}{4}=0.02"

    

  In Excel: =PV(0.02,24,5000,0,1)

       =96461.02